solve by using elimination method
9x+2y+3x=3
2x-y-2z=4
y-12x-12z+6=0

Answer by mananth(16946)

9x+2y+3x=3 There is a typo assumed 3z
2x-y-2z=4
y-12x-12z+6=0
>
>
9x+2y+3z=3 .......1
2x-y-2z=4..........2
y-12x-12z+6=0......3
9 x + 2 y + 3 z = 3
2 x + -1 y -2 z = 4

-12 x + 1 y + -12 z -6



consider equation 1 &2 Eliminate y
Multiply 1 by 1 1
Multiply 2 by 2 1
we get
9 x + 2 y + 3 z = 3
4 x + -2 y + -4 z = 8
Add the two
13 x + 0 y + -1 z = 11 -------------4
consider equation 2 & 3 Eliminate y
Multiply 2 by 1
Multiply 3 by 1
we get
2 x + -1 y + -2 z = 4
-12 x + 1 y + -12 z = -6
Add the two
-10 x + 0 y + -14 z = -2 -------------5
Consider (4) & (5) Eliminate x
Multiply 4 by 10
Multiply (5) by 13
we get 2
130 x + -10 z = 110
-130 x + -182 z = -26
Add the two
0 x + -192 z = 84
/
z = -29 - 7/16

Plug the value of z in (5)
-10 x + 6 1/8 = -2
-10 x = -8 1/8
x = -12 OR 13/16
plug value of x & z in 1
7 5/16 + 2 y + -1 5/16 = 3
2 y = 3 + -7 5/16 + 1 5/16
y= -1.5